A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1850–1859

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

11th [111]. 14th [284]. 1832 [206]. 1842 [171]. 1844 [132]. 1845 [81]. 1849 [82]. 1851 [56, 64]. 1852 [211]. 1853 [171]. 1857 [238]. 1858 [267]. 1859 [284]. 1st [100, 171, 82].

21st [284]. 22nd [250]. 28th [250]. 2nd [56].

30th [175]. 31st [76, 171].

8th [56, 13].

Aberdeen [284]. Abridged [64, 50]. Abstract [14, 233, 54]. Abuses [263]. Academy [14, 14]. Accidents [103, 127].

Accommodation [230]. According [100]. Account [21, 197, 6, 44]. Accounts [133]. Accumulation [41]. Act [206, 211, 81, 132]. Actual [256]. Additional [198, 84]. Addressed [195]. Administration [273, 231, 237, 95]. Admission [223, 256]. Admitted [171]. Adopted [161]. Advancement [56, 185, 250, 284]. Affected [270]. Africans [171]. Ages [225, 236, 163]. Agricultural [139, 124, 4, 242]. Alexander [64]. Also [236, 117, 39, 171]. America [64, 116, 33, 199, 176, 134]. American [192]. Among [198, 196, 49, 112, 270, 42]. Amount [81, 39]. Analysis [169]. Analytical [103, 127]. Anniversary [120, 202, 228, 261, 94, 144, 173]. Annual [12, 37, 15, 224, 264, 66]. Annuities [187]. Anterior [20]. Appendix [117].

Applicability [61]. Appointments [223]. Apprenticeship [226]. April [171, 175]. Area [43]. Arising [150, 181]. Arithmetic [208]. Army [196, 67]. Art [270]. Articles [242, 156]. Ascertain [39]. Assessment [208]. Association [56, 185, 250, 284, 282, 264]. Attempt [68, 39]. Attendance [193, 69]. Auckland [44]. Average [15]. Averages [5].

Bad [142, 215]. Bank [81, 132, 182, 272]. Banking [159, 197]. Barkly [76]. Baynes [82]. Beer [193]. Beer-House [193]. Being [76, 114, 8]. Belfast [107, 178]. Belgium [197]. Benefits [168]. Benoiston [14]. Between [193, 192]. Beyond [224]. Bienfaisance [188, 238]. Bills [122, 39]. Births [157, 236]. Blind [77]. Blisse [52]. Bodies [146]. Bombay [82, 67]. Borough [244]. Boroughs [206, 211]. both [39, 95]. Boys [162]. Bradford [282]. Branches [98]. Brief [272]. Britain [205, 117, 10, 39, 26, 236]. Britanny [14]. British [76, 53, 56, 185, 250, 284, 266, 192, 187, 208, 196, 3, 95, 175, 277, 147]. Brussels [111, 188]. Bruxelles [197]. Building [92, 283]. Buildings [6]. Bullion [233].

C [64]. Calcutta [16]. Capital [41]. Captain [82]. Causes [152, 157, 161]. Census [22, 64, 148, 117, 100, 82, 236, 64]. Centuries [242]. Certain [107, 267]. Changes [81]. Character [257]. Charge [231, 237]. Chateauneuf [14]. Cheltenham [185, 217]. Cheshire [39]. Chest [114]. Chief [38]. Chiefly [95]. Children [69, 61]. China [195]. Chittagon [70]. Cholera [105, 72, 13]. Chronicon [242]. Church [283]. Circulation [81, 132, 39]. Citizens [151]. City [63]. Civil [95]. Class [178, 162]. Classes [69, 168, 270, 27, 28, 41]. Classified [171]. Clearing [169]. Clergy [49]. Close [9]. Clothing [254]. Clyde [92]. Coal [158]. Code [68]. Coffee [62]. Coinage [128].

Colaba [82]. College [281]. Colonial [45]. Colonies [76]. Colony [129, 171, 76]. Coloured [151]. Commerce [25, 126, 175, 140]. Commercial [68]. Commission [83]. Committals [142, 215]. Committee [22]. Communication [107]. Company [54]. Comparative [19, 243, 236]. Compared [205, 273, 224, 38]. Competition [273]. Competitive [223, 256]. Compiled [52]. Compiles [45]. Condition [216, 277]. Conferred [168]. Congrès [188, 238]. Congress [222, 111, 167]. Connected [255]. Connection [192]. Considered [15]. Constitution [182]. Consumption [89]. Contained [107]. Continent [61, 27, 247]. Continental [9]. Contribution [35, 114]. Contributions [3]. Control [224]. Convicts [199]. Copper [55]. Corn [15, 139]. Cost [139, 283, 263]. Cotton [192, 26]. Council [140]. Counties [211]. Countries [43, 205, 111, 39, 206]. Country [33, 271, 132, 27]. Course [281]. Coventry [248]. Credit [197, 245]. Crime [193, 151, 174, 224, 255, 200]. Crimea [196]. Criminal [61, 214]. Crown [203]. Cure [161]. Currant [210]. Currency [81, 132, 177]. Current [139]. Customs [98, 121].

Dated [64]. Deaf [148, 77, 91]. Deaths [122, 157, 224, 236]. Debts [43]. December [64]. Decimal [128]. Deduced [280]. Deduction [200]. Defence [145, 153]. Degradation [151]. Delinquency [161]. Denmark [100]. Department [64]. Departments [14]. Derived [5, 121]. Description [117]. Despatch [76]. Devonshire [13]. Different [273, 5, 41, 236]. Direction [147]. Discharged [54]. Discovery [33]. Disease [106]. Diseases [122, 171, 106, 224, 100, 38]. Dispensaries [248]. Dispensary [248]. Distribution [170]. District [254, 13].

Dockyard [107, 113]. Documents [111]. Dominions [3]. down [45]. drawn [39, 238]. Dumb [148, 77, 91]. Duration [49, 112, 198, 270]. During [122, 14, 169, 15, 7, 157, 145, 153, 206, 211, 19, 257, 175, 44, 280, 105, 25, 139, 39, 283]. Duties [98, 121]. Dwellings [271, 6].

Each [123, 39]. Earlier [208]. Earliest [123]. Earthenware [204]. East [54]. Ecclesiastical [242]. Economic [281]. Economical [223]. Economy [244]. Editor [233]. Edouard [50]. Education [61]. Educational [223]. Effect [142, 215, 75, 71, 98, 156, 140, 230]. Effects [106, 249]. Effort [283]. Eight [175]. Eighteenth [66, 242]. Eighth [250]. Electoral [206, 211, 258, 267]. Elevation [72]. Eleven [171]. embracing [39]. Embroidered [218]. Emigrants [170]. Emigration [7, 83]. Employed [117, 145, 153]. Ended [175]. Engine [92]. England [254, 8, 139, 61, 132, 265, 208, 182, 146, 39, 206, 211, 258, 267, 156, 272, 52]. English [140]. entitled [107]. Enumerated [76]. Epidemic [105]. Equitable [176]. Erected [283, 6]. Especial [83]. Establishment [197]. Estates [83]. Europe [43, 273, 33, 170, 199]. European [38]. Events [224]. Evidence [258]. Examination [256, 26]. Examinations [223]. Exchange [39]. Excise [98, 121]. Exemplified [151, 81, 132, 216]. Existing [192]. Expenditure [34]. Export [10]. Exports [233, 147]. Extent [43, 168, 92, 158, 204, 263]. External [175].

F [56]. F.R.S. [232]. Factory [178]. Farm [61]. Fatal [122]. Fatality [72]. February [267]. Female [194]. Fibres [246]. Field [273]. Fifteen [157]. Fifteenth [94]. Fifth [261]. Financial [266, 187, 277]. Fire [205]. First [56, 281, 144, 145, 153]. Fisheries [115]. Fishery [152]. Five [14, 206, 211, 175]. Fluctuations [15, 157, 224, 39]. Flutuations [152]. Foreign [25, 199, 10, 39]. Foreigners [151, 199]. Former [33, 216]. Founded [178, 52]. Four [233, 95]. Fourteen [122]. Fourth [14]. Fouth [228]. France [15, 25, 160, 236]. Frankfort [238, 63]. Frankfort-on-the-Maine [238, 63]. Free [151, 147, 247, 63]. Freedom [226]. Freehold [104]. French [145, 153, 50]. Friends [278, 262]. Front [2, 31, 60, 87, 110, 137, 166, 191, 221, 253]. Funding [145, 153]. Further [258].

G [64]. General [20, 280, 195]. Geneva [50]. George [280]. Glasgow [161, 216, 209, 204, 243, 249, 283, 162]. Glass [10, 204]. Gold [33, 147]. Good [142, 215]. Government [187, 38]. Governor [76]. Grammar [162]. Granting [187]. Great [263, 205, 117, 10, 39, 26, 236]. Grey [76]. Growth [7]. Guiana [76].

H [64]. Habits [42]. Harbour [216].

Health [265, 178]. Held

[222, 167, 188, 238, 56, 250, 284, 282, 111, 264]. Her [76]. Herring [152]. Historical [129, 123]. History [62, 89, 245, 114]. Hon [76]. Hon. [64]. Honorary [238]. Hospital [171]. Hospitals [168]. House [169, 193]. Hull [105, 106]. Human [224].

Iceland [32]. II [258]. Illegitimacy [280]. Illustrated [246]. Illustrating [139]. Immediate [71, 98]. Importance [194]. Imports [233]. Imposed [28]. Improvement [216]. Improving [133, 6]. Inclusive [75, 157, 39, 38, 95]. Income [88]. Increase [71]. Incumbered [83]. Index [20]. India [54, 53, 3, 34, 95, 175, 229, 277]. Indian [235, 246]. Industrial [244, 163]. Industries [180, 186]. Industry [80]. Inferences [271]. Influence [8, 151, 72, 178, 147]. Influences [223].

Information [1, 24, 30, 48, 59, 79, 86, 102, 109, 131, 136, 155, 165, 184, 190, 213, 220, 241, 252, 276]. infra [233]. Inhabiting [6]. Inland [39]. inquiry [39]. Insane [77]. Instruction [193]. Instructions [14]. Insurance [205]. Intemperate [42]. Interior [64]. International [238, 222, 167, 188]. Introducing [111]. Investigation [256]. Ipswich [56]. Ireland [81, 84, 4, 218, 91]. Irish [83]. Iron [158]. Island [96, 99]. Islands [82].

J. [64]. July [56, 100]. Junior [223]. Justice [95]. Juvenile [161].

Keeping [133]. Kennedy [64]. Kinds [242]. King [281]. Kingdom [55, 7, 180, 186, 128, 28, 257].

Labour [226, 209]. Labourers [271]. Labourers-and [271]. Labouring [6]. Lancashire [39]. Land [104, 208, 84]. Land-Tax [208]. Last [122, 15, 7, 139, 157, 283, 200]. late [76]. Law [123]. Laws [81, 132, 177]. Lawyers [198]. Lead [55]. Learned [198]. Lectures [281]. Leeds [250, 254, 244]. Legislation [273]. Leone [171]. Lepers [77]. Letter [195]. Life [49, 112, 198, 270, 187, 178]. Like [224]. Lists [242]. Literature [270]. Liverpool [225, 271, 264]. Loans [145, 153]. Localities [174]. London [280, 122, 22, 12, 46, 57, 37, 281, 195, 159, 238, 66]. Longevity [198]. Loss [187]. Louis [25]. Lunatics [171]. l’Union [197].

M. [50]. Machine [249]. Machinery [117]. Made [14]. Madras [38]. Magnitude [39]. Maine [238, 63]. Majesty [76]. Making [92, 20]. Mallet [50]. Manchester [225]. Manufacture [254, 192, 10, 204, 218]. March [171, 76]. Marine [92]. Mariners [18]. Maritime [141]. Marriages [157].

Mary [13]. Marylebone [280]. Master [18]. Matter [2, 31, 60, 87, 110, 137, 166, 191, 221, 253]. May [82]. Medical [70, 171, 112, 178, 74]. Meeting [37, 56, 120, 185, 202, 250, 228, 261, 284, 282, 94, 264, 66, 144, 173]. Members [14, 198]. Memoirs [14]. Memorandum [267, 210]. Men [54, 112, 114]. Mental [100]. Mercantile [257]. Methods [145, 153]. Metropolis [21, 157, 51, 231, 237]. Metropolitan [6, 230]. Middle [270]. Military [54, 181]. Mining [180, 186, 160]. Ministers [203]. Miscellanea [97, 101, 108, 119, 125, 130, 135, 143, 154, 164, 172, 179, 183, 189, 201, 207, 212, 219, 259, 268, 274, 286]. Miscellaneous [11, 17, 23, 29, 36, 40, 47, 58, 65, 73, 78, 85, 93]. Miscellenea [149]. MM. [14]. Mobilier [245]. Mode [133]. Money [209, 243]. Money-Rate [209, 243]. Moral [14, 8]. Mortality [122, 105, 150, 181, 196, 18, 27, 42, 74, 38, 67]. Motion [55]. Movement [122, 194]. Mr. [55, 145, 153]. Municipality [283]. Munster [45]. Muslin [218].

Namely [175]. National [43, 133, 264, 163, 282]. Native [199, 38]. Natives [199]. Natural [114]. Nature [168]. Naval [150]. Neutral [140]. Nice [141]. ninth [284]. North [53]. North-West [53]. Northern [115]. Norway [77]. Note [233]. Notes [248, 171, 271, 81, 132, 208, 246, 229]. Notice [217]. Nova [118]. Number [171, 157, 224, 283, 123]. Numbers [5, 163].

Objections [256]. Objects [138]. Observations [280, 5, 198, 279, 84, 114]. Observed [44]. Obtain [117]. October [282, 264]. Office [45, 64]. Officers [196]. Official [45]. Old [122]. One [238, 39]. Operations [150, 181, 182, 272]. Operative [178]. Orders [140]. Origin [13]. Other [205, 224, 9, 242]. Our [126, 163].

Pancras [280, 6]. Paper [278, 267]. Paris [14, 167, 80]. Parish [20, 217]. Parishes [280]. Part [258]. Particular [15, 271]. Particularly [208]. Passed [123, 33]. Passing [81, 132]. Past [277]. Pauper [61]. Pauperism [151]. Pay [203]. Peace [126]. Peculiar [242]. Pendarves [55]. People [8, 163]. Period [171, 7, 208, 123, 95]. Periods [15, 175]. Persons [42]. Philippe [25]. Physical [8, 114]. Pitt [145, 153]. Place [81]. Places [283, 52]. Plate [10]. Plympton [13]. Police [21, 82]. Political [14, 223, 15, 208]. Poor [168, 231, 237, 6]. Poorer [69]. Popular [193]. Population [122, 195, 43, 7, 225, 76, 6, 236]. Porcelain [204]. Portions [267]. Portsea [107, 96, 99, 113]. Portsmouth [107, 113]. Position [146]. pp [233]. Prepared [246]. Present [33, 180, 186, 133, 182, 231, 237, 206, 211, 216, 277, 123, 210]. presented [76]. Preserving [20]. Pretiosum [242]. Prevailing [171]. Prevalence [106]. Preventive [61]. Price [15, 139, 90, 121]. Prices [19, 242, 257, 249]. Principal [161, 156]. Principle [256]. Principles [273]. Prison [142, 215]. Proceedings [22, 46, 57, 239, 285, 188, 238, 264]. Processes [117]. Produce [242, 156]. Produced [71, 3]. Producing [151]. Production [249]. Productions [10]. Productive [80]. Productiveness [98]. Profession [74]. Professions [198]. Professor [281]. Progress [138, 205, 256, 7, 25, 257, 92, 158, 204, 218, 147, 247]. Promotion [282, 264]. Property [88]. Proportion [43, 199]. Proposal [20]. Proposed [161]. Propositions [271]. Prospective [277]. Prospects [266]. Provident [248, 27]. Provinces [53]. Provision [271]. Prussia [9]. Public [223, 256, 265, 34, 229]. Punishment [255]. Purchase [84]. Purpose [111]. Pursuits [270].

Qualifications [223, 163]. Quantity [33]. Quarterly [227, 234, 240, 251, 260, 269, 275, 287].

Race [114]. Railway [103, 127, 230].

Railways [266, 9]. Raised [256, 145, 153]. Rate [231, 237, 42, 74, 209, 243]. Read [267]. Realm [123]. Recorded [123]. Records [45]. Reference [254, 117, 15, 7, 83, 226, 267, 257]. Reform [206, 211]. Reformatories [194]. Reformatory [194, 61, 161]. Registers [20]. Registrar [280, 195]. Registrar-General [280, 195]. Registration [116]. Reign [123, 25]. Relating [236]. Relation [193, 90, 121, 140]. Relative [118, 106, 5, 198]. Relief [231, 237]. Religious [193, 146]. Remarks [54, 194, 10]. Remission [75, 71, 98]. Remote [98]. Report [14, 12, 64, 222, 188, 238, 214, 264]. Reports [240, 257]. Request [238]. Resources [160]. respectively [39]. Restropect [159]. Results [256, 273, 117, 257, 230, 258]. Resume [111, 167]. Return [171, 236]. Returns [280, 55, 227, 234, 251, 260, 269, 275, 287, 194, 117, 171, 214]. Rev. [52]. Revenue [75, 71, 98, 90, 121, 235, 63]. Revenues [43]. Review [43, 272]. Riding [242]. Right [76]. Rise [218]. River [216]. Rivers [216]. Russia [126, 279, 156]. Russian [156, 140].

Same [7]. Sanitary [244, 51]. School [61, 162]. Schools [69, 161]. Science [281, 270, 56, 185, 250, 284, 282, 264]. Sciences [14, 14]. Scotia [118]. Scotland [177, 39, 35, 158, 209, 218, 243]. Season [106]. Second [167, 264, 173]. Secretaries [238]. Secretary [45, 64, 76]. Section [56, 107]. Self [248, 28]. self-Imposed [28]. Self-Supporting [248]. Sent [171]. September [284, 250, 107, 222, 111, 167, 188, 238]. Service [54, 223, 256, 273]. Session [12, 167, 228]. Settlement [208]. Seventeenth [37, 242]. Seventy [15]. Several [199]. Sewing [249]. Shipping [233]. Short [171]. Showing [171, 283]. Sick [171]. Sickness [67]. Sierra [171]. Silver [33]. Since [81, 132]. Sixteenth [12, 242]. Sixth [185]. Size [283, 114]. Slavery [192]. Snaith [242]. Snathense [242]. Social [223, 151, 282, 264]. Societies [104]. Society [22, 12, 46, 57, 37, 278, 270, 41, 238, 6, 66, 120, 138, 202, 239, 228, 285, 261, 262, 76, 94, 144, 173]. Soil [8]. Some [248, 256, 279, 145, 153]. Southwark [280]. Special [254, 157]. Specimens [246]. Spread [13]. St [13]. St. [280, 6]. Standard [265]. State [76, 180, 186, 231, 237, 236]. Statement [19]. Statements [107, 145, 153]. States [199, 9, 64, 116, 151, 170, 134]. Stating [133]. Statistical [217, 129, 43, 222, 271, 21, 111, 167, 146, 6, 123, 44, 22, 12, 46, 57, 37, 120, 138, 202, 239, 228, 285, 261, 107, 94, 238, 66, 144, 173]. Statistics [45, 53, 96, 99, 113, 169, 70, 107, 248, 194, 256, 118, 139, 16, 69, 61, 278, 262, 187, 208, 235, 77, 10, 100, 51, 279, 68, 178, 115, 206, 211, 258, 267, 124, 4, 26, 32, 35, 3, 141, 162, 210, 50, 200, 134, 91, 52, 56, 281]. Stature [114]. Statute [123]. Statutes [123]. Steamboat [92]. Strength [114, 163]. Stuart [64]. Subdivision [8]. Subjects [255]. Substance [76]. Suffolk [174]. Sugar [3]. Suggestions [133]. Summary [270]. Superintendent [82, 64]. Supplementary [278]. Supply [15, 263]. Supporting [248]. supposed [33]. Surface [170]. Sustained [187]. System [116, 61, 226].

T [52]. Table [278, 52]. Tables [233, 54, 117, 171, 236]. Taken [81, 82]. Tax [88, 208]. Taxation [235, 28, 63]. Taxes [75, 71]. Ten [139, 159, 200]. Termed [197].

Terminal [230]. Tested [163]. Tests [223]. Their [114, 238]. them [194]. Theorem [147]. Thereon [231, 237]. Third [202, 282]. Thirteen [171]. Thirty [7, 75]. Thos [232]. Those [271, 98]. Three [198]. Tidal [216]. Time [206, 211, 123, 33, 39]. Times [142, 215]. Tin [55]. Tobacco [89]. Tooke [232, 281]. Touching [271]. Tour [14]. Town [217, 271]. Towns [263]. Trade [55, 10, 26, 257, 158, 210, 147, 247]. Traffic [230]. Translation [50]. Treated [171]. Troops [38]. Twentieth [120]. Twenty [56, 202, 250, 228, 261, 284, 39, 206, 211, 283, 144, 173, 185]. Twenty-Eighth [250]. Twenty-Fifth [261]. Twenty-First [56, 144]. Twenty-Five [206, 211]. Twenty-Fouth [228]. Twenty-ninth [284]. Twenty-second [173]. Twenty-Sixth [185]. Twenty-Third [202]. Two [175].

United [55, 64, 116, 7, 151, 170, 180, 186, 128, 28, 257, 134]. Unity [111]. Universal [68]. Universities [279]. Upon [208, 178, 106, 39, 156]. Upper [270]. Uses [263].

Valuation [84]. Value

[139, 5, 92, 158, 204, 218]. Various

[43, 224, 242, 283]. Victoria [129]. Vienna [222]. View [129, 270, 103, 127, 243, 123, 236]. Villages [176]. Villerme [14]. Vital [70, 16, 278, 262, 32, 35, 50]. Volume [14, 1, 24, 30, 48, 59, 79, 86, 102, 109, 131, 136, 155, 165, 184, 190, 213, 220, 241, 252, 276]. Voluntary [283].

Wages [19, 209, 243, 249]. Wales

[8, 208, 146, 206, 211, 258, 267, 52]. War [126, 145, 153, 156]. was [44]. Washington [64]. Water [263]. Weight [114]. Well [8]. Well-being [8]. West [53, 242, 158, 209, 243]. Whale [115]. Wheat [90, 121]. Will [224]. Within [224, 273, 283, 3]. Woollen [254]. Working [168, 83, 28]. Works [34, 229]. Worship [193, 283, 52].

Year [280, 169, 159, 9, 279, 115, 257, 44]. Years [122, 14, 171, 15, 7, 139, 75, 157, 39, 206, 211, 19, 283, 38, 95, 175, 200]. Yorkshire [242].

Zealand [45, 44, 114]. Zinc [55].

References

Anonymous:1850:VIa

[1] Anonymous. Volume information. Journal of the Statistical Society of London, 13(1):i–iv, February 1850. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338359.

Anonymous:1850:FM

[2] Anonymous. Front matter. Journal of the Statistical Society of London, 13(1): ii, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338360.

Sykes:1850:CSS

[3] W. H. Sykes. Contributions to the statistics of sugar produced within the British Dominions in India. Journal of the Statistical Society of London, 13(1):1–24, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338361.

Porter:1850:ASI

[4] G. R. Porter. Agricultural statistics of Ireland. Journal of the Statistical Society of London, 13(1):25–29, February 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338362.

Guy:1850:RVA

[5] William A. Guy. On the relative value of averages derived from different numbers of observations. Journal of the Statistical Society of London, 13(1):30–45, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338363.

Sykes:1850:SAL

[6] W. H. Sykes. Statistical account of the labouring population inhabiting the buildings at St. Pancras, erected by the Metropolitan Society for Improving the Dwellings of the Poor. Journal of the Statistical Society of London, 13(1):46– 61, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338364.

Danson:1850:PEU

[7] J. T. Danson. On the progress of emigration from the United Kingdom during the last thirty years, with reference to the growth of the population during the same period. Journal of the Statistical Society of London, 13(1):61–63, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338365.

Barton:1850:ISS

[8] John Barton. The influence of subdivision of the soil on the moral and physical well-being of the people of England and Wales. Journal of the Statistical Society of London, 13(1):63–77, February 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338366.

Hebeler:1850:RPO

[9] Bernard Hebeler. Railways in Prussia and other continental states, at the close of the year 1848. Journal of the Statistical Society of London, 13(1):77–80, February 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338367.

Howard:1850:SMP

[10] Henry Howard. Statistics on the manufacture of plate glass in Great Britain from 1760 to 1847, with remarks on foreign productions and the export trade. Journal of the Statistical Society of London, 13(1):80–84, February 1850. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338368.

Anonymous:1850:Ma

[11] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 13(1): 84–96, February 1850. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338369.

Anonymous:1850:SAR

[12] Anonymous. Sixteenth annual report of the Statistical Society of London session 1849–50. Journal of the Statistical Society of London, 13(2):97–102, May 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338376.

MacLaren:1850:OSC

[13] A. C. MacLaren. On the origin and spread of cholera in the 8th District of Plympton St. Mary, Devonshire. Journal of the Statistical Society of London, 13(2):103–134, May 1850. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338377.

Anonymous:1850:ART

[14] Anonymous. Abstract of the report of a tour in the five Departments of Britanny, during the years 1840 and 1841, under instructions from the Academy of the Moral and Political Sciences, made by MM. Benoiston de Chateauneuf and Villerme, members of that Academy; from the fourth volume of the Memoirs of the Academy of the Moral and Political Sciences of Paris. Journal of the Statistical Society of London, 13(2):134– 151, May 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338378.

Danson:1850:FAS

[15] J. T. Danson. On the fluctuations of the annual supply and average price of corn, in France, during the last seventy years, considered with particular reference to the political periods of 1792, 1814, 1830, and 1848. Journal of the Statistical Society of London, 13(2):152– 167, May 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338379.

Finch:1850:VSC

[16] Cuthbert Finch. Vital statistics of Calcutta. Journal of the Statistical Society of London, 13(2):168–182, May 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338380.

Anonymous:1850:Mb

[17] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 13(2): 183–192, May 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338381.

Neison:1850:MMM

[18] F. G. P. Neison. Mortality of master mariners. Journal of the Statistical Society of London, 13(3):193–209, August 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338143.

Porter:1850:CSP

[19] G. R. Porter. On a comparative statement of prices and wages during the years from 1842 to 1849. Journal of the Statistical Society of London, 13(3):210– 218, August 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338144.

Wyatt-Edgell:1850:PPM

[20] E. Wyatt-Edgell. Proposal for preserving and making a general index of the parish registers anterior to 1837. Journal of the Statistical Society of London, 13(3):218–221, August 1850. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338145.

Fletcher:1850:SAP

[21] Joseph Fletcher. Statistical account of the police of the Metropolis. Journal of the Statistical Society of London, 13(3):221–267, August 1850. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338146.

Anonymous:1850:PCC

[22] Anonymous. Proceedings of the census committee of the Statistical Society of London, 1850. Journal of the Statistical Society of London, 13(3):267–269, August 1850. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338147.

Anonymous:1850:Mc

[23] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 13(3): 270–288, August 1850. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338148.

Anonymous:1850:VIb

[24] Anonymous. Volume information. Journal of the Statistical Society of London, 13(4):377–380, December 1850. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338370.

Danson:1850:PFC

[25] J. T. Danson. Progress of the foreign commerce of France during the reign of Louis Philippe, 1830–48. Journal of the Statistical Society of London, 13 (4):289–304, December 1850. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338371.

Porter:1850:ERS

[26] G. R. Porter. Examination of the recent statistics of the cotton trade in Great Britain. Journal of the Statistical Society of London, 13(4):305–312, December 1850. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338372.

Neison:1850:MPC

[27] F. G. P. Neison. Mortality of the provident classes in this country and on the continent. Journal of the Statistical Society of London, 13(4):313–358, December 1850. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338373.

Porter:1850:SIT

[28] G. R. Porter. On the self-imposed taxation of the working classes in the United Kingdom. Journal of the Statistical Society of London, 13(4):358–364, December 1850. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338374.

Anonymous:1850:Md

[29] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 13 (4):364–376, December 1850. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338375.

Anonymous:1851:VIa

[30] Anonymous. Volume information. Journal of the Statistical Society of London, 14(1):i–iv, March 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338107.

Anonymous:1851:FM

[31] Anonymous. Front matter. Journal of the Statistical Society of London, 14(1): ii, March 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338108.

Schleisner:1851:VSI

[32] P. A. Schleisner. Vital statistics of Iceland. Journal of the Statistical Society of London, 14(1):1–10, March 1851. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338109.

Danson:1851:QGS

[33] J. T. Danson. Of the quantity of gold and silver supposed to have passed from America to Europe, from the discovery of the former country to the present time. Journal of the Statistical Society of London, 14(1):11–44, March 1851. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338110.

Sykes:1851:EIP

[34] W. H. Sykes. Expenditure in India on public works from 1837–8 to 1845–6. Journal of the Statistical Society of London, 14(1):45–47, March 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338111.

Stark:1851:CVS

[35] James Stark. Contribution to the vital statistics of Scotland. Journal of the Statistical Society of London, 14(1):48– 87, March 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338112.

Anonymous:1851:Ma

[36] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 14(1): 88–96, March 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338113.

Anonymous:1851:SAM

[37] Anonymous. Seventeenth annual meeting of the Statistical Society of London. Journal of the Statistical Society of London, 14(2):97–108, May 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338237.

Sykes:1851:MCD

[38] W. H. Sykes. Mortality and Chief diseases of the troops under the Madras government European and native, from the years 1842 to 1846 inclusive, compared with the mortality and Chief diseases of 1847. Journal of the Statistical Society of London, 14(2):109–142, May 1851. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338238.

Newmarch:1851:AAM

[39] William Newmarch. An attempt to ascertain the magnitude and fluctuations of the amount of bills of exchange (inland and foreign) in circulation at one time in Great Britain, in England, in Scotland, in Lancashire, and in Cheshire, respectively, during each of the twenty years 1828–1847, both inclusive; and also embracing in the inquiry bills drawn upon foreign countries. Journal of the Statistical Society of London, 14(2):143–183, May 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338239.

Anonymous:1851:Mb

[40] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 14(2): 184–192, May 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338240.

Porter:1851:ACD

[41] G. R. Porter. On the accumulation of capital by the different classes of society. Journal of the Statistical Society of London, 14(3):193–199, September 1851. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338335.

Neison:1851:RMA

[42] F. G. P. Neison. On the rate of mortality among persons of intemperate habits. Journal of the Statistical Society of London, 14(3):200–219, September 1851. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338336.

Brown:1851:SRN

[43] T. J. Brown. A statistical review of the national debts and revenues in proportion to the extent of area and population of the various countries of Europe. Journal of the Statistical Society of London, 14(3):220–226, September 1851. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338337.

Thomson:1851:SAA

[44] Arthur S. Thomson. A statistical account of Auckland, New Zealand, as it was observed during the year 1848. Journal of the Statistical Society of London, 14(3):227–249, September 1851. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338338.

Anonymous:1851:SNM

[45] Anonymous. Statistics of New Munster, New Zealand, down to 1848, compiles from official records in the Colonial Secretary’s Office. Journal of the Statistical Society of London, 14(3):250–261, September 1851. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338339.

Anonymous:1851:PSSa

[46] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 14(3): 261–271, September 1851. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338340.

Anonymous:1851:Mc

[47] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 14 (3):272–288, September 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338341.

Anonymous:1851:VIb

[48] Anonymous. Volume information. Journal of the Statistical Society of London, 14(4):377–380, December 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338251.

Guy:1851:DLA

[49] William A. Guy. On the duration of life among the clergy. Journal of the Statistical Society of London, 14(4):289–297, December 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338252.

Valpy:1851:VSG

[50] Richard Valpy. Vital statistics of Geneva: — an abridged translation from the French of M. Edouard Mallet. Journal of the Statistical Society of London, 14(4):298–337, December 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338253.

Jopling:1851:SSM

[51] R. Thompson Jopling. Sanitary statistics of the Metropolis. Journal of the Statistical Society of London, 14(4):337– 343, December 1851. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338254.

Wyatt-Edgell:1851:SPW

[52] E. Wyatt-Edgell. On the statistics of places of worship in England and Wales, founded on a table compiled by the Rev. T. Blisse. Journal of the Statistical Society of London, 14(4):343–344, December 1851. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338255.

Anonymous:1851:SNW

[53] Anonymous. Statistics of the North-West provinces of British India. Journal of the Statistical Society of London, 14 (4):345–347, December 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338256.

Balfour:1851:RAT

[54] Edward Balfour. Remarks on the abstract tables of the men discharged from the military service of the East India Company. Journal of the Statistical Society of London, 14(4):348–356, December 1851. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338257.

Anonymous:1851:TUK

[55] Anonymous. Trade of the United Kingdom in lead, copper, tin, and zinc. from returns to a motion by Mr. Pendarves. Journal of the Statistical Society of London, 14(4):357–362, December 1851. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338258.

Anonymous:1851:TFM

[56] Anonymous. Twenty-first meeting of the British Association for the Advancement of Science, held at Ipswich, 2nd– 8th July, 1851. Section F. Statistics. Journal of the Statistical Society of London, 14(4):363, December 1851. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338259.

Anonymous:1851:PSSb

[57] Anonymous. Proceedings of the Statistical Society of London. Journal of the Statistical Society of London, 14(4):364, December 1851. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338260.

Anonymous:1851:Md

[58] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 14 (4):364–376, December 1851. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338261.

Anonymous:1852:VIa

[59] Anonymous. Volume information. Journal of the Statistical Society of London, 15(1):i–iv, April 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338307.

Anonymous:1852:FM

[60] Anonymous. Front matter. Journal of the Statistical Society of London, 15(1): ii, April 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338308.

Fletcher:1852:SFS

[61] Joseph Fletcher. Statistics of the farm school system of the continent, and of its applicability to the preventive and reformatory education of pauper and criminal children in England. Journal of the Statistical Society of London, 15(1):1– 49, April 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338309.

Crawford:1852:HC

[62] John Crawford. History of coffee. Journal of the Statistical Society of London, 15(1):50–58, April 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338310.

Sykes:1852:TRF

[63] W. H. Sykes. Taxation and revenue of the Free City of Frankfort-on-the-Maine. Journal of the Statistical Society of London, 15(1):59–63, April 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338311.

Anonymous:1852:CUS

[64] Anonymous. Census of the United States of America, abridged from the report of J. C. G. Kennedy, the Superintendent of the census, to the Hon. Alexander H. H. Stuart, Secretary of the Department of the Interior: Dated Census Office, Washington, December 1, 1851. Journal of the Statistical Society of London, 15(1):64–84, April 1852. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338312.

Anonymous:1852:Ma

[65] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 15(1): 85–96, April 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338313.

Tooke:1852:EAM

[66] T. Tooke. Eighteenth annual meeting of the Statistical Society of London. Journal of the Statistical Society of London, 15(2):97–99, June 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338299.

Sykes:1852:MSB

[67] W. H. Sykes. Mortality and sickness of the Bombay Army, 1848–9. Journal of the Statistical Society of London, 15(2): 100–107, June 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338300.

Levi:1852:CSA

[68] Leone Levi. On commercial statistics, and an attempt at a universal commercial code. Journal of the Statistical Society of London, 15(2):108–114, June 1852. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338301.

Fletcher:1852:SAS

[69] Joseph Fletcher. Statistics of the attendance in schools for children of the poorer classes. Journal of the Statistical Society of London, 15(2):115–117, June 1852. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338302.

Bedford:1852:VMS

[70] J. R. Bedford. On the vital and medical statistics of Chittagon. Journal of the Statistical Society of London, 15(2):117– 150, June 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338303.

Guy:1852:IEP

[71] Dr. Guy. On the immediate effect produced on the revenue by the remission or increase of taxes. Journal of the Statistical Society of London, 15(2):150– 154, June 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338304.

Farr:1852:IEF

[72] William Farr. Influence of elevation on the fatality of cholera. Journal of the Statistical Society of London, 15(2):155– 183, June 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338305.

Anonymous:1852:Mb

[73] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 15(2): 184–192, June 1852. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338306.

Neison:1852:RMM

[74] F. G. P. Neison. On the rate of mortality in the medical profession. Journal of the Statistical Society of London, 15 (3):193–222, September 1852. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338126.

Guy:1852:ERT

[75] Dr. Guy. On the effect of the remission of taxes on the revenue in the thirty years from 1822 to 1851 inclusive. Journal of the Statistical Society of London, 15 (3):223–227, September 1852. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338127.

Grey:1852:PCB

[76] Earl Grey. On the population of the Colony of British Guiana, as enumerated on the 31st of March, 1851. being the substance of a despatch from Governor Barkly, presented to the Society of the Right Hon. Earl Grey, Her Majesty’s late Secretary of State for the Colonies. Journal of the Statistical Society of London, 15(3):228–249, September 1852. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338128.

Holst:1852:SIB

[77] Professor Holst and A. S. Oliver Massey. On the statistics of the insane, blind, deaf and dumb, and lepers, of Norway. Journal of the Statistical Society of London, 15(3):250–256, September 1852. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338129.

Anonymous:1852:Mc

[78] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 15 (3):256–288, September 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338130.

Anonymous:1852:VIb

[79] Anonymous. Volume information. Journal of the Statistical Society of London, 15(4):377–380, December 1852. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979627.

Porter:1852:PIP

[80] G. R. Porter. On the productive industry of Paris. Journal of the Statistical Society of London, 15(4):289–307, December 1852. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979628.

Gilbart:1852:LCI

[81] J. W. Gilbart. On the laws of the currency in Ireland, as exemplified in the changes that have taken place in the amount of bank notes in circulation in Ireland, since the passing of the Act of 1845. Journal of the Statistical Society of London, 15(4):307–326, December 1852. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979629.

Sykes:1852:CIB

[82] Colonel Sykes. On the census of the islands of Bombay and Colaba, taken on the 1st of May, 1849, by Captain Baynes, Superintendent of Police. Journal of the Statistical Society of London, 15(4):327–338, December 1852. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2979630.

Locke:1852:IEE

[83] John Locke. On Irish emigration, with especial reference to the working of the Incumbered Estates Commission. Journal of the Statistical Society of London, 15(4):339–345, December 1852. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2979631.

Locke:1852:AOV

[84] John Locke. Additional observations on the valuation and purchase of land in Ireland. Journal of the Statistical Society of London, 15(4):345–350, December 1852. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979632.

Anonymous:1852:Md

[85] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 15 (4):350–376, December 1852. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2979633.

Anonymous:1853:VIa

[86] Anonymous. Volume information. Journal of the Statistical Society of London, 16(1):i–iv, March 1853. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338099.

Anonymous:1853:FM

[87] Anonymous. Front matter. Journal of the Statistical Society of London, 16(1): ii, March 1853. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338100.

Farr:1853:IPT

[88] William Farr. The income and property tax. Journal of the Statistical Society of London, 16(1):1–44, March 1853. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338101.

Crawford:1853:HCT

[89] John Crawford. On the history and consumption of tobacco. Journal of the Statistical Society of London, 16(1):45– 52, March 1853. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338102.

Guy:1853:RPW

[90] Dr. Guy. On the relation of the price of wheat to the revenue. Journal of the Statistical Society of London, 16(1):53– 68, March 1853. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338103.

Wilde:1853:SDD

[91] W. R. Wilde. Statistics of the deaf and dumb in Ireland. Journal of the Statistical Society of London, 16(1):69– 77, March 1853. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338104.

Strang:1853:PEV

[92] John Strang. On the progress, extent, and value of steamboat building and marine engine making on the Clyde. Journal of the Statistical Society of London, 16(1):77–84, March 1853. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338105.

Anonymous:1853:Ma

[93] Anonymous. Miscellaneous. Journal of the Statistical Society of London, 16(1): 85–96, March 1853. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338106.

Overstone:1853:FAM

[94] Lord Overstone. Fifteenth anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 16(2):97–102, June 1853. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338215.

Sykes:1853:ACJ

[95] Colonel Sykes. Administration of civil justice in British India, for a period of four years, chiefly from 1845 to 1848, both inclusive. Journal of the Statistical Society of London, 16(2):103–136, June 1853. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338216.

Anonymous:1853:SIPa

[96] Anonymous. Statistics of the Island of Portsea. Journal of the Statistical Society of London, 16(2):137–168, June 1853. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338217.

Anonymous:1853:Mb

[97] Anonymous. Miscellanea. Journal of the Statistical Society of London, 16(2):169– 192, June 1853. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338218.

Guy:1853:IRE

[98] Dr. Guy. On the immediate and remote effect of the remission of customs and excise duties on the productiveness of those branches of the revenue. Journal of the Statistical Society of London, 16 (3):193–200, September 1853. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338114.

Anonymous:1853:SIPb

[99] Anonymous. Statistics of the Island of Portsea. Journal of the Statistical Society of London, 16(3):201–243, September 1853. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338115.

Hubertz:1853:SMD

[100] J. R. Hubertz. Statistics of mental diseases in Denmark, according to the census of July 1st, 1847. Journal of the Statistical Society of London, 16(3):244– 274, September 1853. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338116.

Anonymous:1853:Mc

[101] Anonymous. Miscellanea. Journal of the Statistical Society of London, 16(3): 275–288, September 1853. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338117.

Anonymous:1853:VIb

[102] Anonymous. Volume information. Journal of the Statistical Society of London, 16(4):373–384, December 1853. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338131.

Neison:1853:AVR

[103] F. G. P. Neison. Analytical view of railway accidents. Journal of the Statistical Society of London, 16(4):289–337, December 1853. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338132.

Beggs:1853:FLS

[104] Thomas Beggs. Freehold land societies. Journal of the Statistical Society of London, 16(4):338–346, December 1853. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338133.

Cooper:1853:CMH

[105] Henry Cooper. On the cholera mortality in Hull during the epidemic of 1849. Journal of the Statistical Society of London, 16(4):347–351, December 1853. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338134.

Cooper:1853:RPD

[106] Henry Cooper. On the relative prevalence of diseases in Hull, and the effects of season upon disease. Journal of the Statistical Society of London, 16 (4):352–355, December 1853. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338135.

Bentham:1853:CSC

[107] Lady Bentham. On certain statements contained in a communication to the Statistical Section at Belfast, in September, 1852, entitled “Statistics of Portsea and Portsmouth Dockyard”. Journal of the Statistical Society of London, 16(4):356–362, December 1853. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338136.

Anonymous:1853:Md

[108] Anonymous. Miscellanea. Journal of the Statistical Society of London, 16(4):

363–372, December 1853. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338137.

Anonymous:1854:VIa

[109] Anonymous. Volume information. Journal of the Statistical Society of London, 17(1):i–iv, March 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338348.

Anonymous:1854:FM

[110] Anonymous. Front matter. Journal of the Statistical Society of London, 17(1): ii, March 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338349.

Levi:1854:RSC

[111] Leone Levi. Resume of the statistical Congress, held at Brussels, September 11th, 1853, for the purpose of introducing unity in the statistical documents of all countries. Journal of the Statistical Society of London, 17(1):1–14, March 1854. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338350.

Guy:1854:DLA

[112] William A. Guy. On the duration of life among medical men. Journal of the Statistical Society of London, 17(1):15– 23, March 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338351.

Anonymous:1854:SPP

[113] Anonymous. Statistics of Portsea and Portsmouth Dockyard. Journal of the Statistical Society of London, 17(1):24– 26, March 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338352.

Thomson:1854:CNH

[114] A. S. Thomson. Contribution to the natural history of the New Zealand race of men; being observations on their stature, weight, size of chest, and physical strength. Journal of the Statistical Society of London, 17(1):27–33, March 1854. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338353.

Munroe:1854:SNW

[115] Henry Munroe. Statistics of the northern whale fisheries, from the year 1772 to 1852. Journal of the Statistical Society of London, 17(1):34–42, March 1854. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338354.

Curtis:1854:SRU

[116] Josiah Curtis. On the system of registration in the United States of America. Journal of the Statistical Society of London, 17(1):43–44, March 1854. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338355.

Cheshire:1854:RCG

[117] Edward Cheshire. The results of the census of Great Britain in 1851, with a description of the machinery and processes employed to obtain the returns; also an appendix of tables of reference. Journal of the Statistical Society of London, 17(1):45–72, March 1854. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338356.

Cheshire:1854:SRN

[118] Edward Cheshire. Statistics relative to Nova Scotia in 1851. Journal of the Statistical Society of London, 17(1):73– 80, March 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338357.

Anonymous:1854:Ma

[119] Anonymous. Miscellanea. Journal of the Statistical Society of London, 17(1):81– 96, March 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338358.

Anonymous:1854:TAM

[120] Anonymous. Twentieth anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 17(2):97–102, June 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338269.

Guy:1854:RPW

[121] Dr. Guy. On the relation of the price of wheat to the revenue derived from customs and excise duties. Journal of the Statistical Society of London, 17(2):103– 116, June 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338270.

Angus:1854:ONB

[122] John Angus. Old and new bills of mortality; movement of the population; deaths and fatal diseases in London during the last fourteen years. Journal of the Statistical Society of London, 17(2): 117–142, June 1854. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338271.

Tayler:1854:SHV

[123] William Tayler. A statistical and historical view of the statute law of the realm, and of the number of statutes passed in each reign from the earliest recorded period to the present time. Journal of the Statistical Society of London, 17(2):143– 158, June 1854. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338272.

Paull:1854:AS

[124] Samuel Paull. On agricultural statistics. Journal of the Statistical Society of London, 17(2):159–167, June 1854. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338273.

Anonymous:1854:Mb

[125] Anonymous. Miscellanea. Journal of the Statistical Society of London, 17(2):168– 192, June 1854. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338274.

Danson:1854:OCR

[126] J. T. Danson. Our commerce with Russia, in peace and war. Journal of the Statistical Society of London, 17(3):193– 218, September 1854. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338319.

Neison:1854:AVR

[127] F. G. P. Neison. Analytical view of railway accidents. Journal of the Statistical Society of London, 17(3):219–242, September 1854. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338320.

Minasi:1854:DCU

[128] Frederic James Minasi. On a decimal coinage for the United Kingdom. Journal of the Statistical Society of London, 17(3):243–258, September 1854. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338321.

Bell:1854:HSV

[129] G. M. Bell. Historical and statistical view of the colony of Victoria. Journal of the Statistical Society of London, 17 (3):259–274, September 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338322.

Anonymous:1854:Mc

[130] Anonymous. Miscellanea. Journal of the Statistical Society of London, 17(3): 275–288, September 1854. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338323.

Anonymous:1854:VIb

[131] Anonymous. Volume information. Journal of the Statistical Society of London, 17(4):373–388, December 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338288.

Gilbart:1854:LCE

[132] J. W. Gilbart. The laws of the currency, as exemplified in the circulation of country bank notes in England, since the passing of the Act of 1844. Journal of the Statistical Society of London, 17 (4):289–321, December 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338289.

Jellicoe:1854:SIP

[133] Charles Jellicoe. Suggestions for improving the present mode of keeping and stating the national accounts. Journal of the Statistical Society of London, 17 (4):322–325, December 1854. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338290.

Welton:1854:SUS

[134] Thomas Abercrombie Welton. Statistics of the United States of America. Journal of the Statistical Society of London, 17 (4):326–357, December 1854. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338291.

Anonymous:1854:Md

[135] Anonymous. Miscellanea. Journal of the Statistical Society of London, 17(4): 358–372, December 1854. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338292.

Anonymous:1855:VIa

[136] Anonymous. Volume information. Journal of the Statistical Society of London, 18(1):i–iv, March 1855. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338118.

Anonymous:1855:FM

[137] Anonymous. Front matter. Journal of the Statistical Society of London, 18(1): ii, March 1855. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338119.

Anonymous:1855:OPS

[138] Anonymous. Objects and progress of the Statistical Society. Journal of the Statistical Society of London, 18(1):1– 2, March 1855. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338120.

Danson:1855:CPC

[139] J. T. Danson. On the current price, and the cost price, of corn, in England, during the last ten years, as illustrating the value of agricultural statistics.

Journal of the Statistical Society of London, 18(1):3–20, March 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338121.

Waddilove:1855:ERO

[140] Alfred Waddilove. The effect of the recent orders in Council, in relation to English, Russian, and neutral commerce. Journal of the Statistical Society of London, 18(1):21–33, March 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338122.

Sykes:1855:SNM

[141] Colonel Sykes. Statistics of nice maritime. Journal of the Statistical Society of London, 18(1):34–73, March 1855. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338123.

Clay:1855:EGB

[142] Rev. John Clay. On the effect of good or bad times on committals to prison. Journal of the Statistical Society of London, 18(1):74–79, March 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338124.

Anonymous:1855:Ma

[143] Anonymous. Miscellanea. Journal of the Statistical Society of London, 18(1):80– 96, March 1855. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338125.

Tooke:1855:TFA

[144] Thomas Tooke. Twenty-first anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 18(2):97–103, June 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338201.

Newmarch:1855:LRMa

[145] William Newmarch. On the loans raised by Mr. Pitt during the First French War, 1793–1801; with some statements in defence of the methods of funding employed. Journal of the Statistical Society of London, 18(2):104–140, June 1855. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338202.

Mann:1855:SPR

[146] Horace Mann. On the statistical position of religious bodies in England and Wales. Journal of the Statistical Society of London, 18(2):141–159, June 1855. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338203.

Valpy:1855:PDB

[147] Richard Valpy. The progress and direction of British exports, and the influence theorem of free trade and gold. Journal of the Statistical Society of London, 18(2):160–173, June 1855. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338204.

Buxton:1855:CDD

[148] David Buxton. The census of the deaf and dumb in 1851. Journal of the Statistical Society of London, 18(2):174–185, June 1855. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338205.

Anonymous:1855:Mb

[149] Anonymous. Miscellenea. Journal of the Statistical Society of London, 18(2):186– 200, June 1855. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338206.

Hodge:1855:MAN

[150] William Barwick Hodge. On the mortality arising from Naval operations. Journal of the Statistical Society of London, 18(3):201–221, September 1855. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338314.

Everest:1855:ISD

[151] Robert Everest. On the influence of social degradation in producing pauperism and crime, as exemplified in the free coloured citizens and foreigners in the United States. Journal of the Statistical Society of London, 18(3):222–239, September 1855. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338315.

Cleghorn:1855:CFH

[152] John Cleghorn. On the causes of the flutuations in the Herring fishery. Journal of the Statistical Society of London, 18 (3):240–242, September 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338316.

Newmarch:1855:LRMb

[153] William Newmarch. On the loans raised by Mr. Pitt during the First French War, 1793–1801; with some statements in defence of the methods of funding employed. Journal of the Statistical Society of London, 18(3):242–284, September 1855. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338317.

Anonymous:1855:Mc

[154] Anonymous. Miscellanea. Journal of the Statistical Society of London, 18(3): 285–300, September 1855. CODEN ????

ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338318.

Anonymous:1855:VIb

[155] Anonymous. Volume information. Journal of the Statistical Society of London, 18(4):393–407, December 1855. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338278.

Valpy:1855:EWR

[156] Richard Valpy. The effect of the War, in Russia and England, upon the principal articles of Russian produce. Journal of the Statistical Society of London, 18 (4):301–311, December 1855. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338279.

Guy:1855:FNB

[157] William A. Guy. On the fluctuations in the number of births, deaths, and marriages, and in the number of deaths from special causes, in the Metropolis, during the last fifteen years, from 1840 to 1854, inclusive. Journal of the Statistical Society of London, 18(4):312–329, December 1855. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338280.

Strang:1855:PEV

[158] John Strang. On the progress, extent, and value of the coal and iron trade of the West of Scotland. Journal of the Statistical Society of London, 18(4): 330–332, December 1855. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338281.

Gilbart:1855:TYR

[159] James William Gilbart. A ten years’ restropect of London banking. Journal of the Statistical Society of London, 18 (4):333–344, December 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338282.

Lack:1855:MRF

[160] H. Reader Lack. The mining resources of France, 1841 to 1852. Journal of the Statistical Society of London, 18(4): 345–355, December 1855. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338283.

MCallum:1855:JDP

[161] A. K. M’Callum. Juvenile delinquency– its principal causes and proposed cure, as adopted in the Glasgow reformatory schools. Journal of the Statistical Society of London, 18(4):356–363, December 1855. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338284.

Tennent:1855:SGG

[162] Andrew Tennent. Statistics of a Glasgow grammar school class of 115 Boys. Journal of the Statistical Society of London, 18(4):364–366, December 1855. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338285.

Yeats:1855:ONS

[163] John Yeats. On our national strength; tested by the numbers, the ages, and the industrial qualifications of the people. Journal of the Statistical Society of London, 18(4):367–377, December 1855. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338286.

Anonymous:1855:Md

[164] Anonymous. Miscellanea. Journal of the Statistical Society of London, 18(4):

378–392, December 1855. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338287.

Anonymous:1856:VIa

[165] Anonymous. Volume information. Journal of the Statistical Society of London, 19(1):i–iv, March 1856. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338168.

Anonymous:1856:FM

[166] Anonymous. Front matter. Journal of the Statistical Society of London, 19(1): ii, March 1856. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338169.

Levi:1856:RSS

[167] Leone Levi. Resume of the second session of the International Statistical Congress held at Paris, September, 1855. Journal of the Statistical Society of London, 19(1):1–11, March 1856. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338170.

Guy:1856:NEB

[168] William A. Guy. On the nature and extent of the benefits conferred by hospitals on the working classes and the poor. Journal of the Statistical Society of London, 19(1):12–27, March 1856. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338171.

Babbage:1856:ASC

[169] Charles Babbage. Analysis of the statistics of the clearing house during the year 1839. Journal of the Statistical Society of London, 19(1):28–48, March 1856. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338172.

Everest:1856:DEE

[170] Robert Everest. On the distribution of the emigrants from Europe over the surface of the United States. Journal of the Statistical Society of London, 19(1):49– 59, March 1856. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338173.

Clarke:1856:SNP

[171] Robert Clarke. Short notes of the prevailing diseases in the colony of Sierra Leone, with a return of the sick Africans sent to hospital in eleven years, and classified medical returns for the years 1853– 4. Also tables showing the number of lunatics admitted to hospital in a period of thirteen years, and the number treated from 1st April, 1842, to 31st March, 1853. Journal of the Statistical Society of London, 19(1):60–81, March 1856. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338174.

Anonymous:1856:Ma

[172] Anonymous. Miscellanea. Journal of the Statistical Society of London, 19(1):81– 96, March 1856. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338175.

Vane:1856:TSA

[173] Harry George Vane. Twenty-second anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 19(2):97–101, June 1856. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338262.

Glyde:1856:LCS

[174] Mr. John Glyde. Localities of crime in Suffolk. Journal of the Statistical Society of London, 19(2):102–106, June 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338263.

Sykes:1856:ECB

[175] Colonel Sykes. The external commerce of British India during two periods of years; namely, the eight years (ended 30th April) 1834–35 to 1841–42, and the five years 1849–50 to 1853–54. Journal of the Statistical Society of London, 19(2):107–126, June 1856. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338264.

Pare:1856:EVA

[176] William Pare. Equitable villages in America. Journal of the Statistical Society of London, 19(2):127–143, June 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338265.

Gilbart:1856:LCS

[177] J. W. Gilbart. The laws of the currency in Scotland. Journal of the Statistical Society of London, 19(2):144–169, June 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338266.

Malcolm:1856:IFL

[178] A. G. Malcolm. The influence of factory life on the health of the operative, as founded upon the medical statistics of this class at Belfast. Journal of the Statistical Society of London, 19(2):170– 181, June 1856. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338267.

Anonymous:1856:Mb

[179] Anonymous. Miscellanea. Journal of the Statistical Society of London, 19(2):182– 200, June 1856. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338268.

Hunt:1856:PSMa

[180] Robert Hunt. The present state of the mining industries of the United Kingdom. Journal of the Statistical Society of London, 19(3):201–218, September 1856. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338188.

Hodge:1856:MAM

[181] William Barwick Hodge. On the mortality arising from military operations. Journal of the Statistical Society of London, 19(3):219–271, September 1856. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338189.

Jellicoe:1856:BEP

[182] Charles Jellicoe. The Bank of England; its present constitution and operations. Journal of the Statistical Society of London, 19(3):272–283, September 1856. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338190.

Anonymous:1856:Mc

[183] Anonymous. Miscellanea. Journal of the Statistical Society of London, 19(3): 284–304, September 1856. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338191.

Anonymous:1856:VIb

[184] Anonymous. Volume information. Journal of the Statistical Society of London, 19(4):401–408, December 1856. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338293.

Anonymous:1856:BAA

[185] Anonymous. British Association for the Advancement of Science. Twenty-sixth meeting, 1856. (Cheltenham). Journal of the Statistical Society of London, 19 (4):305–310, December 1856. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338294.

Hunt:1856:PSMb

[186] Robert Hunt. The present state of the mining industries of the United Kingdom. Journal of the Statistical Society of London, 19(4):311–324, December 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338295.

Hendriks:1856:FSB

[187] Frederick Hendriks. On the financial statistics of British government life annuities (1808–1855), and on the loss sustained by government in granting annuities. Journal of the Statistical Society of London, 19(4):325–384, December 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338296.

Lumley:1856:RPI

[188] W. G. Lumley. Report on the proceedings of the International Congrès de Bienfaisance, held at Brussels, September, 1856. Journal of the Statistical Society of London, 19(4):385–389, December 1856. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338297.

Anonymous:1856:Md

[189] Anonymous. Miscellanea. Journal of the Statistical Society of London, 19(4): 390–400, December 1856. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338298.

Anonymous:1857:VIa

[190] Anonymous. Volume information. Journal of the Statistical Society of London, 20(1):i–iv, March 1857. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338156.

Anonymous:1857:FM

[191] Anonymous. Front matter. Journal of the Statistical Society of London, 20(1): ii, March 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338157.

Danson:1857:ECB

[192] J. T. Danson. On the existing connection between American slavery and the British cotton manufacture. Journal of the Statistical Society of London, 20(1): 1–21, March 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338158.

Clay:1857:RBC

[193] John Clay. On the relation between crime, popular instruction, attendance on religious worship, and beer-house. Journal of the Statistical Society of London, 20(1):22–32, March 1857. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338159.

Carpenter:1857:ISR

[194] Mary Carpenter. On the importance of statistics to the reformatory movement, with returns from female reformatories, and remarks on them. Journal of the Statistical Society of London, 20(1):33– 40, March 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338160.

Bowring:1857:PCL

[195] John Bowring. The population of China. a letter addressed to the Registrar-General London. Journal of the Statistical Society of London, 20(1):41–53, March 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338161.

Jopling:1857:MAO

[196] R. Thompson Jopling. On the mortality among officers of the British Army in the Crimea. Journal of the Statistical Society of London, 20(1):54–60, March 1857. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338162.

Lumley:1857:ABE

[197] William Golden Lumley. An account of the banking establishment in Belgium, termed l’Union du Credit de Bruxelles. Journal of the Statistical Society of London, 20(1):61–65, March 1857. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338163.

Guy:1857:DLA

[198] William A. Guy. On the duration of life among lawyers; with additional observations on the relative longevity of the members of the three learned professions. Journal of the Statistical Society of London, 20(1):65–71, March 1857. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338164.

Everest:1857:PFN

[199] R. Everest. On the proportion of foreigners to natives, and of foreign and native convicts, in several states of Europe and America. Journal of the Statistical Society of London, 20(1):72–76, March 1857. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338165.

Walsh:1857:DSC

[200] Richard Hussey Walsh. A deduction from the statistics of crime for the last ten years. Journal of the Statistical Society of London, 20(1):77–78, March 1857. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338166.

Anonymous:1857:Ma

[201] Anonymous. Miscellanea. Journal of the Statistical Society of London, 20(1):79– 96, March 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338167.

Anonymous:1857:TTA

[202] Anonymous. Twenty-third anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 20(2):97–101, June 1857. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338342.

Farr:1857:PMC

[203] William Farr. On the pay of Ministers of the Crown. Journal of the Statistical Society of London, 20(2):102–131, June 1857. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338343.

Strang:1857:PEV

[204] John Strang. On the progress, extent, and value of the porcelain, earthenware, and glass manufacture of Glasgow. Journal of the Statistical Society of London, 20(2):132–134, June 1857. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338344.

Brown:1857:PFI

[205] Samuel Brown. On the progress of fire insurance in Great Britain, as compared with other countries. Journal of the Statistical Society of London, 20(2):135– 168, June 1857. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338345.

Newmarch:1857:ESCa

[206] William Newmarch. On the electoral statistics of the countries and boroughs in England and Wales during the twentyfive years from the Reform Act of 1832 to the present time. Journal of the Statistical Society of London, 20(2):169– 234, June 1857. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338346.

Anonymous:1857:Mb

[207] Anonymous. Miscellanea. Journal of the Statistical Society of London, 20(2):235– 240, June 1857. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338347.

Hendriks:1857:SBL

[208] Frederick Hendriks. On the statistics of the British land-tax assessment, and particularly of England and Wales, from 1636 to 1856; with notes upon the political arithmetic of the earlier period of its settlement. Journal of the Statistical Society of London, 20(3):241–307, September 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338138.

Strang:1857:MRW

[209] John Strang. On the money-rate of wages of labour in Glasgow and the West of Scotland. Journal of the Statistical Society of London, 20(3):308–312, September 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338139.

Travers:1857:MPS

[210] John Ingram Travers. Memorandum on the present statistics of the currant trade. Journal of the Statistical Society of London, 20(3):313, September 1857. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338140.

Newmarch:1857:ESCb

[211] William Newmarch. On the electoral statistics of the counties and boroughs in England and Wales during the twentyfive years from the Reform Act of 1852 to the present time. Journal of the Statistical Society of London, 20(3):314–340, September 1857. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338141.

Anonymous:1857:Mc

[212] Anonymous. Miscellanea. Journal of the Statistical Society of London, 20(3): 341–364, September 1857. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338142.

Anonymous:1857:VIb

[213] Anonymous. Volume information. Journal of the Statistical Society of London, 20(4):457–467, December 1857. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338230.

Tartt:1857:RCR

[214] W. M. Tartt. Report on criminal returns. Journal of the Statistical Society of London, 20(4):365–377, December 1857. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338231.

Clay:1857:EGB

[215] John Clay. On the effect of good or bad times on committals to prison. Journal of the Statistical Society of London, 20 (4):378–388, December 1857. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338232.

Strang:1857:ITR

[216] John Strang. The improvement of tidal rivers, as exemplified by the former and present condition of the river and harbour of Glasgow. Journal of the Statistical Society of London, 20(4):389–396, December 1857. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338233.

Beamish:1857:SNT

[217] Richard Beamish. Statistical notice of the town and Parish of Cheltenham. Journal of the Statistical Society of London, 20(4):397–423, December 1857. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338234.

Strang:1857:RPV

[218] John Strang. On the rise, progress, and value of the embroidered muslin manufacture of Scotland and Ireland. Journal of the Statistical Society of London, 20 (4):424–428, December 1857. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338235.

Anonymous:1857:Md

[219] Anonymous. Miscellanea. Journal of the Statistical Society of London, 20(4): 429–456, December 1857. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338236.

Anonymous:1858:VIa

[220] Anonymous. Volume information. Journal of the Statistical Society of London, 21(1):i–iv, March 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338207.

Anonymous:1858:FM

[221] Anonymous. Front matter. Journal of the Statistical Society of London, 21(1): ii, March 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338208.

Brown:1858:RIS

[222] Samuel Brown. Report on the International Statistical Congress, held at Vienna, September, 1857. Journal of the Statistical Society of London, 21(1):1– 17, March 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338209.

Chadwick:1858:ESE

[223] Edwin Chadwick. On the economical, social, educational, and political influences of competitive examinations, as tests of qualifications for admission to the junior appointments in the public service. Journal of the Statistical Society of London, 21(1):18–51, March 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338210.

Guy:1858:AFN

[224] William A. Guy. On the annual fluctuations in the number of deaths from various diseases, compared with like fluctuations in crime, and in other events within and beyond the control of the human will. Journal of the Statistical Society of London, 21(1):52–86, March 1858. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338211.

Danson:1858:APL

[225] J. T. Danson. On the ages of the population in Liverpool and Manchester. Journal of the Statistical Society of London, 21(1):87–93, March 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338212.

Napier:1858:ASR

[226] James Robert Napier. The apprenticeship system, in reference to the freedom of labour. Journal of the Statistical Society of London, 21(1):94–97, March 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338213.

Anonymous:1858:QRa

[227] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 21(1):98–114, March 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338214.

Anonymous:1858:TFA

[228] Anonymous. Twenty-fouth anniversary meeting of the Statistical Society. Session 1857–58. Journal of the Statistical Society of London, 21(2):115–120, June 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338149.

Sykes:1858:NPW

[229] Colonel Sykes. Notes on public works in India. Journal of the Statistical Society of London, 21(2):121–155, June 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338150.

Wilkinson:1858:MRT

[230] William Arthur Wilkinson. Metropolitan Railway terminal accommodation, and its effect on traffic results. Journal of the Statistical Society of London, 21(2):156–168, June 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338151.

Lumley:1858:PSAa

[231] W. G. Lumley. On the present state of the administration of the relief to the poor in the Metropolis, and the charge of the poor rate thereon. Journal of the Statistical Society of London, 21(2):169– 197, June 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338152.

Anonymous:1858:TTF

[232] Anonymous. Thos Tooke F.R.S. Journal of the Statistical Society of London, 21(2):198–201, June 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338153.

Anonymous:1858:NEF

[233] Anonymous. Note by Editor on the four abstract tables of imports, exports, shipping, and bullion, at pp. 214–217 infra. Journal of the Statistical Society of London, 21(2):201, June 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338154.

Anonymous:1858:QRb

[234] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 21(2):202–222, June 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338155.

Hendriks:1858:SIR

[235] Frederick Hendriks. On the statistics of Indian revenue and taxation. Journal of the Statistical Society of London, 21 (3):223–296, September 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338224.

Willich:1858:TRS

[236] Charles M. Willich and E. T. Scargill. Tables relating to the state of the population of Great Britain at the census of 1851, with a comparative view, at the different ages, of the population of France; also a comparative return of births and deaths, 1838–1854. Journal of the Statistical Society of London, 21 (3):297–307, September 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338225.

Lumley:1858:PSAb

[237] W. G. Lumley. On the present state of the administration of relief to the poor in the Metropolis, and charge of the poor rate thereon. Journal of the Statistical Society of London, 21(3):308–338, September 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338226.

Roberts:1858:RPC

[238] Henry Roberts. Report on the proceedings of the Congrès International de Bienfaisance, held in Frankfort-on-the-Maine, September, 1857. Drawn up for the Statistical Society of London, at the request of one of their Honorary Secretaries. Journal of the Statistical Society of London, 21(3):339–344, September 1858. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338227.

Anonymous:1858:PSS

[239] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 21(3):345–346, September 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338228.

Anonymous:1858:QRc

[240] Anonymous. Quarterly reports. Journal of the Statistical Society of London, 21 (3):347–368, September 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338229.

Anonymous:1858:VIb

[241] Anonymous. Volume information. Journal of the Statistical Society of London, 21(4):491–502, December 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338324.

Robinson:1858:CPS

[242] Charles Best Robinson. Chronicon Pretiosum Snathense; or lists of prices of various kinds of agricultural produce, and of other articles, in the Ecclesiastical Peculiar of Snaith, in the West Riding of Yorkshire, in the Sixteenth, Seventeenth, and Eighteenth Centuries. Journal of the Statistical Society of London, 21(4):369–420, December 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338325.

Strang:1858:CVM

[243] John Strang. Comparative view of the money-rate of wages in Glasgow and the West of Scotland, in 1851, 1856, and 1858. Journal of the Statistical Society of London, 21(4):421–426, December 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338326.

Baker:1858:ISE

[244] Robert Baker. On the industrial and sanitary economy of the Borough of Leeds, in 1858. Journal of the Statistical Society of London, 21(4):427–443, December 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338327.

Newmarch:1858:RHC

[245] William Newmarch. On the recent history of the credit mobilier. Journal of the Statistical Society of London, 21 (4):444–453, December 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338328.

Sadler:1858:NIF

[246] J. H. Sadler and Colonel Sykes. Notes on Indian fibres illustrated by prepared specimens. Journal of the Statistical Society of London, 21(4):454–456, December 1858. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338329.

vandeMaeren:1858:PFT

[247] M. Corr van de Maeren. On the progress of free trade on the continent. Journal of the Statistical Society of London, 21 (4):457–459, December 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338330.

Bracebridge:1858:NSS

[248] Charles H. Bracebridge. Notes on self-supporting dispensaries, with some statistics of the Coventry Provident Dispensary. Journal of the Statistical Society of London, 21(4):460–463, December 1858. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338331.

Strang:1858:SMG

[249] John Strang. The sewing machine in Glasgow, and its effects on production, prices, and wages. Journal of the Statistical Society of London, 21(4):464–467, December 1858. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338332.

Anonymous:1858:TEM

[250] Anonymous. Twenty-eighth meeting of the British Association for the Advancement of Science, held at Leeds, 22nd– 28th September, 1858. Journal of the Statistical Society of London, 21(4):468– 469, December 1858. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338333.

Anonymous:1858:QRd

[251] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 21 (4):470–490, December 1858. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338334.

Anonymous:1859:VIa

[252] Anonymous. Volume information. Journal of the Statistical Society of London, 22(1):i–vi, March 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338192.

Anonymous:1859:FM

[253] Anonymous. Front matter. Journal of the Statistical Society of London, 22(1): ii, March 1859. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338193.

Baines:1859:WME

[254] Edward Baines. On the woollen manufacture of England; with special reference to the Leeds clothing district. Journal of the Statistical Society of London, 22(1):1–34, March 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338194.

Tartt:1859:SCC

[255] W. M. Tartt. On subjects connected with crime and punishment. Journal of the Statistical Society of London, 22(1): 35–43, March 1859. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338195.

Chadwick:1859:PPC

[256] Edwin Chadwick. On the progress of the principle of competitive examination for admission into the public service, with statistics of actual results and an investigation of some of the objections raised. Journal of the Statistical Society of London, 22(1):44–75, March 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338196.

Section:1859:MRC

[257] Editorial Section. Mercantile reports of the character and results of the trade of the United Kingdom during the year 1858; with a reference to the progress of prices, 1851–9. Journal of the Statistical Society of London, 22(1):76–100, March 1859. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338197.

Newmarch:1859:ESEa

[258] William Newmarch. On the electoral statistics of England and Wales, 1856–8.

— Part II. Results of further evidence. Journal of the Statistical Society of London, 22(1):101–168, March 1859. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338198.

Anonymous:1859:Ma

[259] Anonymous. Miscellanea. Journal of the Statistical Society of London, 22(1):169– 177, March 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338199.

Anonymous:1859:QRa

[260] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 22(1):178–202, March 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338200.

Anonymous:1859:TFA

[261] Anonymous. Twenty-fifth anniversary meeting of the Statistical Society. Journal of the Statistical Society of London, 22(2):203–207, June 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338179.

Fox:1859:VSS

[262] Joseph John Fox. On the vital statistics of the Society of Friends. Journal of the Statistical Society of London, 22(2):208– 231, June 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338180. See supplement

[278].

Strang:1859:WSG

[263] John Strang. On water supply to great towns: Its extent, cost, uses, and abuses. Journal of the Statistical Society of London, 22(2):232–249, June 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338181.

Tayler:1859:RPS

[264] William Tayler. Report of the proceedings of the second annual meeting of The National Association for The Promotion of Social Science, held at Liverpool, in October, 1858. Journal of the Statistical Society of London, 22(2):250– 252, June 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338182.

Greenhow:1859:SPH

[265] E. Headlam Greenhow. On a standard of public health for England. Journal of the Statistical Society of London, 22(2): 253–270, June 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338183.

Brown:1859:FPB

[266] Samuel Brown. On the financial prospects of British railways. Journal of the Statistical Society of London, 22(2): 271–296, June 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338184.

Newmarch:1859:ESEb

[267] William Newmarch. On the electoral statistics of England and Wales, 1856–8 — memorandum with reference to certain portions of the paper read in February, 1858. Journal of the Statistical Society of London, 22(2):297–305, June 1859. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338185.

Anonymous:1859:Mb

[268] Anonymous. Miscellanea. Journal of the Statistical Society of London, 22(2):306– 313, June 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338186.

Anonymous:1859:QRb

[269] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 22(2):314–335, June 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338187.

Guy:1859:DLA

[270] William A. Guy. On the duration of life as affected by the pursuits of literature, science, and art: with a summary view of the duration of life among the upper and middle classes of society. Journal of the Statistical Society of London, 22 (3):337–361, September 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338382.

Danson:1859:PIS

[271] J. T. Danson. Propositions and inferences, with statistical notes, touching the provision of country dwellings for town labourers-and in particular for those of the town of Liverpool. Journal of the Statistical Society of London, 22 (3):362–377, September 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338383.

Valpy:1859:BRO

[272] Richard Valpy. A brief review of the operations of the Bank of England in 1857. Journal of the Statistical Society of London, 22(3):378–380, September 1859. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338384.

Chadwick:1859:RDP

[273] Edwin Chadwick. Results of different principles of legislation and administration in Europe; of competition for the field, as compared with competition within the field, of service. Journal of the Statistical Society of London, 22 (3):381–420, September 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338385.

Anonymous:1859:Mc

[274] Anonymous. Miscellanea. Journal of the Statistical Society of London, 22(3): 421–430, September 1859. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338386.

Anonymous:1859:QRc

[275] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 22 (3):431–454, September 1859. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338387.

Anonymous:1859:VIb

[276] Anonymous. Volume information. Journal of the Statistical Society of London, 22(4):563–578, December 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338622.

Sykes:1859:PPP

[277] Colonel Sykes. The past, present, and prospective financial condition of British India. Journal of the Statistical Society of London, 22(4):455–480, December 1859. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338623.

Fox:1859:STP

[278] J. J. Fox. Supplementary table to paper “On the Vital Statistics of the Society of Friends”. Journal of the Statistical Society of London, 22(4):481–483, December 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338624. See [262].

Koolomzine:1859:SOS

[279] A. Koolomzine. Some observations and statistics on the universities of Russia in the year 1856. Journal of the Statistical Society of London, 22(4):484–490, December 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338625.

Acton:1859:OIL

[280] William Acton. Observations on illegitimacy in the London parishes of St. Marylebone, St. Pancras, and St. George’s, Southwark, during the year 1857; deduced from the returns of the Registrar-General. Journal of the Statistical Society of London, 22(4):491–505, December 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338626.

Anonymous:1859:FCL

[281] Anonymous. First course of lectures by the Tooke Professor of Economic Science and Statistics, in King’s College, London. Journal of the Statistical Society of London, 22(4):506, December 1859. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338627.

Heywood:1859:TMN

[282] James Heywood. Third meeting of the National Association for the Promotion of Social Science, held at Bradford, in October, 1859. Journal of the Statistical Society of London, 22(4):507–510, December 1859. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338628.

Strang:1859:CBG

[283] John Strang. Church building in Glasgow, showing the number, size, and cost of the various places of worship erected within the municipality during the last twenty years, 1839–59, through voluntary effort. Journal of the Statistical Society of London, 22(4):511–515, December 1859. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338629.

Anonymous:1859:TNM

[284] Anonymous. Twenty-ninth meeting of the British Association for the Advancement of Science, held at Aberdeen, 14th– 21st September, 1859. Journal of the Statistical Society of London, 22(4):516– 517, December 1859. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338630.

Anonymous:1859:PSS

[285] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 22(4):518–520, December 1859. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338631.

Anonymous:1859:Md

[286] Anonymous. Miscellanea. Journal of the Statistical Society of London, 22(4): 521–540, December 1859. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338632.

Anonymous:1859:QRd

[287] Anonymous. Quarterly returns. Journal of the Statistical Society of London, 22 (4):541–562, December 1859. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338633.